### v2

# Beyond RMSE and MAE: Introducing EAUC to unmask hidden bias and unfairness in dyadic regression models

Abstract
Dyadic regression models, which predict real-valued outcomes for pairs of entities, are fundamental in many domains (e.g. predicting the rating of a user to a product in Recommender Systems) and promising and under exploration in others (e.g. approximating the adequate dosage of a drug for a patient in personalized pharmacology). In this work, we prove that non-uniformity in the observed value distributions of individual entities leads to severely biased predictions in state-of-the-art models, skewing predictions towards the average of observed past values for the entity and providing worse-than-random predictive power in eccentric yet equally important cases. We show that global error metrics like Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE) are insufficient to capture this phenomenon, which we name eccentricity bias, and we introduce Eccentricity-Area Under the Curve (EAUC) as a novel complementary metric that can quantify it in all studied models and datasets. We also show the adequateness of EAUC, demonstrating with naive de-biasing corrections that a lower model bias correlates with a lower EAUC and vice-versa. This work contributes a bias-aware evaluation of dyadic regression models to prevent unfairness and risks in critical real-world applications of such systems.
Index Terms:
Regression on Dyadic Data, Fairness, Ethics in Artificial Intelligence, Machine Learning
I Introduction
Within the rapidly evolving fields of Artificial Intelligence (AI) and Machine Learning, dyadic data systems, which operate on data involving pairs of entities such as users and items, have emerged as a pivotal component and wield significant influence over decisions made by individuals or that may affect them. This research paper delves into three interrelated aspects: regression over dyadic data, the evaluation of such tasks, and the pervasive issue of unfairness biases in AI.
Dyadic data systems play a significant role in our data-driven world, being at the core of recommendation engines, personalized content delivery, and countless applications which involve understanding complex relationships between entities like products, movies, or even potential job candidates. In these contexts, regression over dyadic data becomes the process of predicting values for a given pair of entities, such as user ratings for specific products or evaluating the suitability of a job applicant. These predictions can influence anything, from purchasing decisions to employment opportunities.
However, within these critical tasks, the presence of biases related to unfairness can have profound implications, such as disparate impacts on minority or vulnerable groups, unequal access to opportunities, and discriminatory decision-making processes [1] [2]. From a legal perspective, regulations and guidelines are emerging globally to ensure fairness and ethics in AI systems. For instance, the European Union’s AI Act will regulate that AI systems must not be biased such they “perpetuate historical patterns of discrimination” [3], and the US government has emphasized that agencies and offices must act to “address algorithmic discrimination” [4]. Such regulations underscore the need for fairness in algorithmic decision-making to prevent discriminatory outcomes.
In the particular context of regression over dyadic data or dyadic regression, unfairness biases may manifest in real-world contexts as suboptimal recommendations that prioritize highly-rated items over more adequate ones for a user’s needs, or amplify harmful messages in social media, favouring the creation of toxic echo chambers[5][6].
One of the core challenges in evaluating regression tasks over dyadic data is the multi-dimensionality of such data. Typically, entities in the data are categorized into two sets and referred to as “users” and “items”; we adopt this notation for the rest of this work. Because each observed value belongs to a particular pair of entities, each entity (user or item) can have a distinct local distribution of observed values. In this work, we prove such an additional layer of local value distributions in the data renders global error metrics like Root Mean Square Error (RMSE) or Mean Absolute Error (MAE) insufficient to assess the fairness aspect of dyadic regression models. We show that when these distributions are not uniform, RMSE-driven evaluation unfairly favours predictions biased towards the average past observed values of each user and item. This results in models that minimize RMSE but perform poorly for non-regular cases, with errors surpassing those of a random predictor.
Figure 1: Graphical summary of the covered topics and contributions of this research work, including the definition of eccentricity bias in dyadic regression, the limitations of global error metrics like RMSE or MAE for comprehensively evaluating said tasks, and the proposal of EAUC as a novel metric that can quantify the degree of eccentricity bias (and its subsequent unfairness) a dyadic regression model suffers.
In this research, we explore essential aspects of regression over dyadic data, the limitations of benchmarks guided by global error metrics like RMSE, and the necessity of mitigating underlying unfairness in existing state-of-the-art models. Our contributions in the field of regression over dyadic data are summarized in Figure 1 and include:
  • Showing that state-of-the-art models exhibit a strong predictive bias, skewing towards the average observed values of each user and item’s past interactions, leading to algorithmic unfairness and usage risks.
  • Demonstrating how global error metrics like RMSE and MAE, typically the only gauges of performance in dyadic regression, are insufficient to detect such a predictive bias.
  • Proposing a novel metric EAUC (Eccentricity-Area Under the Curve) which, complementing classic metrics like RMSE or MAE, provides a bias-aware assessment of model performance.
All our contributions are reported based on findings made on five top state-of-the-art models and six benchmark datasets for dyadic regression. Furthermore, we demonstrate the correlation of a lower EAUC with smaller eccentricity bias and vice-versa by using naive de-biasing corrections and find that the severity of the bias caused by a given dataset can be quantified with an entity-wise adaptation of the Kolmogorov-Smirnov uniformity statistic.
II Background and Motivation
In this section, we introduce readers to dyadic data and dyadic regression -the data and task contexts of interest-, overview the common evaluation policies for these systems and their historical foundation, and unmask hidden biases in dyadic regression models that point to limitations in said evaluation protocols.
II-A Defining Dyadic Data and Dyadic Regression
Dyadic data, in its essence, consists of interactions between pairs of entities, where each observation involves a user interacting with an item. A dyadic data task typically involves performing a prediction given a pair of entities (e.g. predicting whether a job applicant and a vacant position are compatible) or predicting the most optimal pair of entities given the requisites of the task (e.g. predicting the most suitable applicant for a job position).
Formally, let $\mathcal{U}$ and $\mathcal{I}$ denote the set of users and items, respectively. A dyad can be represented as $(u,i)$, where $u\in\mathcal{U}$ and $i\in\mathcal{I}$ are a particular user and item. These dyads constitute the dataset $\mathcal{D}=\{(u,i)\mid u\in U,i\in I\}$.
This work’s focus lies on regression tasks conducted over dyadic data, where the objective is to predict a real value for a given dyadic observation by determining the optimal regressor function $f$:
$f:\mathcal{U}\times\mathcal{I}\rightarrow\mathbb{R}$
such that $f$ predicts the most adequate $\hat{r}_{ui}\in\mathbb{R}$ for the pair $(u,i)$.
In the State of the Art, dyadic regression tasks most commonly appear within Recommender Systems (RS), where the goal is to present relevant items to users, and dyadic regression can be used to model Learning to Rank (L2R) problems, like ranking the top-N movies for a given user. In L2R, the precise magnitude of the value $r_{ui}$ assigned by $f$ to a dyad $(u,i)$ is not necessarily relevant, as it is only used to define the ranking, which is the final output of the system. However, RS is a thriving field and dyadic regression is commonly used as a surrogate learning goal; as such, some L2R benchmark datasets and models are in practice the State of the Art for dyadic regression and have been of particular interest in this research.
Beyond RS, regression on dyadic data finds applications in various critical domains like finance, to know the most profitable market price of a commodity for different buyers or countries [7], or biomedicine, to predict the adequate dosage of a drug required for an individual’s treatment [8][9]. In these, unlike in L2R tasks, the precise magnitude of $\hat{r}_{ui}$ is crucial. As such, dyadic regression can affect people or their decisions, and predictions should be accurate, unbiased and fair due to their potential real-world impact.
II-B State-of-the-Art Evaluation: Netflix Prize Competition and RMSE
For more than one decade, Root Mean Squared Error (RMSE) has stood as the prevailing evaluation metric for dyadic regression tasks, especially in the field of Recommender Systems. Its adoption can be attributed to seminal events, notably the Netflix Prize [10], a competition initiated by Netflix in 2006 that challenged researchers outperform the company’s existing movie recommendation system, Cinematch. This competition provided a fertile ground for innovation [11] and brought significant attention to the field [12][13].
RMSE was the official means of assessing algorithmic performance during the Netflix Prize competition, and is in essence a global error metric, formally defined as:
$RMSE=\sqrt{\frac{1}{N}\sum_{k=1}^{N}(r_{k}-\hat{r_{k}})^{2}}$ (1)
where $N$ is the total number of predictions, $r_{i}$ is the observed value, and $\hat{r_{i}}$ is the model’s prediction. In the Netflix Prize, observed values were discrete ratings $r_{ui}\in{1,2,3,4,5}$, and models predicted real-valued ratings $\hat{r}_{ui}\in[1,5]$.
RMSE gained prominence due to its intuitive interpretation (lower RMSE values indicating higher prediction accuracy), and the reliance on RMSE and other global error metrics like Mean Absolute Error (MAE) has continued to influence the field way after the Netflix Prize. Similar popular datasets like MovieLens [14] further cemented global error metrics as RMSE or MAE as the sole standard for evaluating regression tasks over dyadic data [15][16].
However, in a dyadic data context, RMSE alone may neglect issues beyond global prediction error, like the existence of biases or knowledge generalization concerns; factors such as the sparsity of past observations or distinct observed value distributions of each user or item could obscure the evaluation.
III Motivation: Limitations of RMSE in Dyadic regression tasks
In this section, we delve into the limitations of global error metrics like RMSE or MAE as the sole means for evaluating performance in dyadic regression tasks. As traditionally used, they evaluate considering the dataset as a whole without distinguishing between different users, items, user-item pairs or their associated traits. While suitable for classic regression tasks, this global error measurement may not sufficiently capture the intricacies of dyadic data.
To exemplify the limitations of RMSE we turn to a practical example using the Netflix Prize data [10] and a Matrix Factorization (MF) model [17], a benchmark dataset and model from the State of the Art; both are further described in Sections V-A and V-B, respectively.
Table I shows the Test RMSE for the samples of each unique rating (1 to 5) and the entire Test set. Although the system performs better for reviews with ratings 3 or 4, the most common in the data ($\sim 28\%$ and $\sim 33\%$ of ratings, respectively), predictions are always notably better than random, suggesting the absence of any important biases.
TABLE I: RMSE of an MF model, seggregated by rating
Rating 1 Rating 2 Rating 3 Rating 4 Rating 5 All Ratings
MF 1.681 1.082 0.636 0.594 0.914 0.822
Random 2.309 1.528 1.154 1.528 2.311 1.697
  • Values obtained by an MF model in the Netflix Prize dataset. The rating is the value $r_{i}$ to be predicted, i.e. a user’s score to a given movie.
However, dyadic data may exhibit different rating distributions in different users or items. For instance, in the case of movie ratings, users may be less or more critical and movies may be less or more critically acclaimed. As such, this apparent lack of bias requires further examination.
Let us define the average observed value $\bar{r}_{u}$ for a user $u$ or $\bar{r}_{i}$ for an item $i$ as:
$\bar{r}_{u}=\frac{1}{|\mathcal{I}_{u}|}\sum_{i\in\mathcal{I}_{u}}r_{ui}\quad\bar{r}_{i}=\frac{1}{|\mathcal{U}_{i}|}\sum_{u\in\mathcal{U}_{i}}r_{ui}$ (2)
where $I_{u}$ is the set of items that $u$ interacted with, $U_{i}$ is the set of users who interacted with item $i$, and $r_{ui}$ is an observed value.
Additionally, given a $(u,i)$ pair, let the average of their $\bar{r}_{u}$ and $\bar{r}_{i}$ be their Dyadic Mean Value ($DMV_{ui}$). Formally:
$DMV_{ui}=\frac{\bar{r_{u}}+\bar{r_{i}}}{2}$ (3)
Finally, let the eccentricity $Ecc_{ui}$ of an observed value $r_{ui}$ measure of how much $r_{ui}$ deviates from the expected value of a $(u,i)$ interaction based on $u$’s and $i$’s past observations, as:
$Ecc_{ui}=|r_{ui}-DMV_{ui}|$ (4)
Using these concepts, Figure 2 displays the relationship between sample eccentricity $Ecc_{ui}$ and prediction error $|\hat{r}_{ui}-r_{ui}|$ of the model and a uniform random predictor. We observe that, when the $r_{ui}$ to be predicted is close to the expected value $DMV_{ui}$ of the interaction (samples with low $Ecc_{ui}$), the model’s predictions are very precise. However, the accuracy of predictions significantly deteriorates as $r_{ui}$ is further away from $DMV_{ui}$ (samples with high $Ecc_{ui}$) Worryingly, at high $Ecc_{ui}$, MF’s predictions are considerably worse than predicting at random.
Figure 2: Eccentricity $Ecc_{ui}$ vs. prediction error $|\hat{r}_{ui}-r_{ui}|$ of an MF model and a uniform random predictor in the Netflix dataset; mean and standard deviation of error over 5 runs is provided. The $y=x$ line corresponds to predicting always the $DMV_{ui}$ of each sample.
To observe this condition from an alternative perspective, Figure 3 shows the relation between ratings and predictions for three different subsets of users and items. The left-most chart depicts all test cases, the central chart depicts test cases $(u,i)$ where $2.0\leq DMV_{ui}\leq 2.5$ (user-item combinations with low average observed values), and the right-most chart features test cases with $4.0\leq DMV_{ui}\leq 4.5$ (user-item combinations with high average observed values). MF is mostly unbiased if viewing the dataset as a whole, but in the low and high $DMV_{ui}$ specific scenarios we observe that the predictions are severely skewed towards the $DMV_{ui}$ of the dyad: the model assumes a critical user will not ever like a low-rated film even if they gave it a high rating, and vice-versa: we refer to this phenomenon as eccentricity bias. As such, the knowledge generalization of the model is lower than an RMSE-guided evaluation would suggest: most of the global error minimization is achieved by blindly skewing predictions toward the average observed past values, rather than by learning latent user and item characteristics.
Depending on the application, this bias can constitute a source of unfairness and a potential risk of use. For instance, if predicting the adequate drug dosage of a specific pharmacological compound $i$which is typically given in large doses- for a given patient $u$who has a high resistance to treatments-, a prediction always close to $DMV_{ui}$ may result in severe overdosing of the patient and dangerous side effects; ultimately, patient safety would be compromised.
Figure 3: Average and standard deviation of Test predicted vs. observed values for an MF model in the Netflix Prize dataset, in three different scenarios: all user-item pairs, and user-item pairs with low and high average ratings (low and high $DMV_{ui}$).
IV Bias-aware Evaluation of Dyadic Regression Models
This section introduces EAUC, an alternative metric that can detect and quantify eccentricity bias and its subsequent unfairness in dyadic regression models’ predictions by providing a holistic view of their performance: a good model should excel not only in predicting near the $DMV_{ui}$ of the dyad, i.e. the expected value of the (u,i) interaction, but also in cases with higher eccentricity, where the model must prove it has understood $u$’s and $i$’s latent characteristics rather than memorizing $\bar{r}_{u}$ and $\bar{r}_{i}$. Such eccentric cases are equally essential, often representing critical edge cases of tasks where an AI system would be most useful.
IV-A Eccentricity-Area Under the Curve (EAUC)
The underlying principle of EAUC is to, by accounting for the eccentricity of samples, ensure that the measured performance of a model can reflect its ability to handle cases that deviate significantly from the average past values of the entities involved. This overcomes the main limitation of global error metrics like RMSE or MAE, where the performance in these uncommon cases has a negligible effect on the measured error.
EAUC can be computed through the following process, also codified in Algorithm 1:
  1. 1.
    Calculate in the Train set the mean observed values $\bar{r}_{u}$ and $\bar{r}_{i}$ for each user $u$ or item $i$.
  2. 2.
    For each test sample $j:(u,i,r_{ui},\hat{r}_{ui})$, calculate:
    • $Ecc_{j}$, the eccentricity of the observed value $r_{ui}$.
    • $\epsilon_{j}$, the absolute prediction error $|\hat{r}_{ui}-r_{ui}|$.
  3. 3.
    Sort Test samples in ascending order of eccentricity, creating an ordered sequence of pairs $(Ecc_{j},\epsilon_{j})$ with increasing $Ecc_{j}$ values.
  4. 4.
    Integrate to determine the area $A$ under the curve defined by eccentricities $Ecc$ and absolute errors $\epsilon$.
  5. 5.
    Normalize the area $A$ within the range [0, 1] by dividing by the area of a rectangle with sides $Ecc_{max}$ and $\epsilon_{max}$ (both can be computed as $max(r_{ui})-min(r_{ui})$).
Algorithm 1 EAUC (Eccentricity-Area Under the Curve)
Inputs:
  • Train samples $D_{\text{train}}=\{(u,i,r_{ui})\}$
  • Test samples $D_{\text{test}}=\{(u,i,r_{ui}$
  • ($u\in\mathcal{U},i\in\mathcal{I},r_{ui}\in\mathbb{R},\hat{r}_{ui}\in\mathbb{R}$)
Outputs: Calculated score EAUC
Initialize $\bar{r}_{users}\leftarrow\mathbb{R}^{|\mathcal{U}|}$, $\bar{r}_{items}\leftarrow\mathbb{R}^{|\mathcal{I}|}$
for $j=1$ to $|\mathcal{U}|$ do
5:     $\bar{r}_{users_{j}}\leftarrow\frac{1}{|D_{\text{train}_{j}}|}\sum_{(u,i,r_{ui})\in D_{\text{train}_{j}}}r_{ui}$
end for
for $j=1$ to $|\mathcal{I}|$ do
     $\bar{r}_{items_{j}}\leftarrow\frac{1}{|D_{\text{train}_{j}}|}\sum_{(u,i,r_{ui})\in D_{\text{train}_{j}}}r_{ui}$
end for
10:Initialize $Ecc\leftarrow\mathbb{R}^{|\mathcal{D}_{\text{test}}|}$ and $\epsilon\leftarrow\mathbb{R}^{|\mathcal{D}_{\text{test}}|}$
for $j=1$ to $|D_{\text{test}}|$ do
     $(u,i,r_{ui},\hat{r}_{ui})\leftarrow D_{\text{test}_{j}}$
     $Ecc_{j}\leftarrow|r_{ui}-\frac{1}{2}\cdot(\bar{r}_{users_{u}}+\bar{r}_{items_{i}})|$
     $\epsilon_{j}\leftarrow|\hat{r}_{ui}-r_{ui}|$
15:end for
Sort $(Ecc_{j},\epsilon_{j})$ pairs by ascending $Ecc_{j}$
Initialize $A\leftarrow 0$
for $j=2$ to $|D_{\text{test}}|$ do
     $\Delta A_{j}\leftarrow\frac{1}{2}\cdot(Ecc_{j}-Ecc_{j-1})\cdot(\epsilon_{j}+\epsilon_{j-1})$
20:     $A\leftarrow A+\Delta A_{j}$
end for
EAUC $\leftarrow\frac{A}{(max(r_{ui})-min(r_{ui}))^{2}};(u,i,r_{ui},\hat{r}_{ui})\in D_{\text{test}}$
IV-B Interpretation and usage of EAUC
EAUC produces a value in the range [0, 1], where lower values indicate lower eccentricity bias and better predictive power. An ideal model with perfect predictions not biased towards $DMV_{ui}$ would achieve an EAUC of 0, and a naive predictor that always predicts the $DMV_{ui}$ would achieve an EAUC of 0.5, as long as it is tested in samples with eccentricities that span the range $[0,max(r_{ui})]$.
EAUC offers a valuable complement to traditional metrics like RMSE and MAE, which quantify average errors but are not aware of the predictive bias learnt by the model to optimize its predictions. By integrating EAUC into the evaluation framework, we extend the evaluation by quantifying the predictive biases towards average observed values of users and items. Since an eccentricity bias could potentially lead to unfairness, EAUC also behaves as a measure of fairness in dyadic regression tasks.
IV-C Measuring Dataset Difficulty for Eccentricity Bias
Using the same dyadic regression dataset, EAUC can be used to measure and compare the eccentricity bias of different models. Between datasets, however, a difference in EAUC could be caused not only by model adequateness, but also due to the characteristics or difficulty of the dataset. Dataset difficulty metrics are used in Machine Learning to assess the upfront difficulty of tasks or datasets and contextualize model performance [18, 19, 20]. In dyadic data, dataset sparsity and scarcity are used to measure local and global quantities of information, but do not characterize the local distributions of observed values of individual entities, which may also influence the learning of models.
We could argue that, as state-of-the-art models learn by minimizing a global error metric like MSE, their eccentricity bias can be more pronounced if the observed values for each entity are concentrated around its average rating, Imagine a user $u$’s observed values follow a normal distribution $X\sim\mathcal{N}(\bar{r}_{u},\,\sigma^{2})$, i.e. they are normally distributed with mean $\bar{r}_{u}$ and standard deviation $\sigma$. The lower $\sigma$ is, the more a model will be rewarded for biasing predictions towards $\bar{r}_{u}$ instead of learning latent characteristics that would provide generalisable knowledge. Conversely, if $u$’s observed values are uniformly distributed, the model must learn these characteristics to minimize prediction errors.
Based on this reasoning, we propose adapting the Kolmogorov-Smirnov uniformity statistic $D_{KS}$ to approximate the severity of the eccentricity biases induced by a given dataset in models learnt with global error optimization; the underlying idea is to assess the uniformity of average observed value distributions for each entity.
For a given dataset $\mathcal{D}$, we compute this entity-wise uniformity statistic, which we term $D_{KS_{\mathcal{D}}}$, as:
$D_{KS_{\mathcal{D}}}=\frac{\sum_{u\in|\mathcal{U}|}D_{KS_{u}}+\sum_{i\in|\mathcal{I}|}D_{KS_{i}}}{|\mathcal{U}|+|\mathcal{I}|}$ (5)
where $D_{KS_{u}}$ or $D_{KS_{i}}$ is the Kolmogorov-Smirnov statistic of a given user’s or item’s distribution of observed values, measuring its maximum deviation to a uniform distribution $X\sim\mathcal{U}(min(r_{ui}),max(r_{ui}))$. A lower $D_{KS_{\mathcal{D}}}$ signifies more uniform observed value distributions across the dataset’s users and items, where models are less tempted to prioritize the $DMV_{ui}$ for any $(u,i)$ pair. Conversely, a higher $D_{KS_{\mathcal{D}}}$ indicates non-uniform distributions, luring models into biasing predictions towards the respective $DMV_{ui}$ and succumbing to eccentricity bias.
V Experimental Setup
In this section, we cover different aspects of our experimental design, including datasets, models and other implementation details, with the objective of analyzing the extent of the studied bias and evaluating the adequateness of EAUC and the difficulty of RMSE to detect such a bias.
V-A Datasets
We focus on dyadic regression datasets where each observation involves entities of two different types and an associated numerical value to be predicted. Table II shows basic statistics of the selected datasets, including the $D_{KS_{\mathcal{D}}}$ measurement of dataset difficulty. This selection provides a variety of data sources, context tasks, sample quantity and sparsity, and uniformity of local value distributions.
TABLE II: Basic statistics of the used datasets.
Dataset Observations Density Users Items $\mathbf{D_{KS_{\mathcal{D}}}}$
Movielens 100K [^1] https://grouplens.org/datasets/movielens/ [14] 100,000 6.30% 944 1,683 0.415
Movielens 1Mfootnotemark: [14] 1,000,209 4.50% 6,040 3,706 0.412
Movielens 10Mfootnotemark: [14] 10,000,054 1.30% 69,878 10,677 0.437
Netflix Prize [^2] https://archive.org/download/nf_prize_dataset.tar [10] 100,480,507 1.20% 480,189 17,770 0.429
Douban Monti [^3] https://github.com/fmonti/mgcnn [21] 136,066 1.50% 3,000 3,000 0.466
Tripadvisor [22] 491,170 0.20% 25,388 12,384 0.539
  • The density ($|\mathcal{D}_{\text{train}}|/(|\mathcal{U}|\times|\mathcal{I}|)$) indicates the quantity of implicit feedback information available per entity (user or item).
Each of these datasets offers a scenario where the goal is to predict the numerical value a user would assign to a particular item (e.g., movies, TV series, restaurants). We apply a 90/10 random train/test data split, a common partitioning strategy that exposes models to the cold start problem [23]. As most users in the original Tripadvisor data only have 1 sample, we only keep those users and items with $\geq 10$ observed interactions. To maintain a strict dyadic regression protocol, no extra information per sample is available other than user and item identifiers.
V-B Models
This work employs a diverse array of state-of-the-art dyadic regression models to assess their underlying eccentricity biases and the behaviour of EAUC in measuring them. The models used in this study are Bayesian SVD++ [24] (matrix decomposition), Glocal-K [25] (auto-encoders and convolution kernels), GC-MC [26] (graph auto-encoders), and MF [17] (simple matrix factorization). We also use two naive predictors as baselines: a random uniform predictor (Random) and a predictor which always outputs the $DMV_{ui}$ of the sample (Dyad Average). This model selection covers different algorithmic approaches used to obtain baseline and state-of-the-art performances in dyadic regression, allowing for a comprehensive evaluation of the prevalence of eccentricity bias, the limitations of global error metrics like RMSE and MAE, and the ability of EAUC to quantify its presence or absence.
V-C Experimental Implementation Details
For each model and dataset, we acquire results using the hyperparameters for that dataset described in the model’s original publication, to respect result reproducibility. If such configurations cannot be found in the article, we fall back on those defined in the official code provided by the authors. The full relation of used hyperparameters for each model and dataset can be found in Appendix A.
Model training and evaluation is done on a dedicated machine with 16GB RAM, an NVIDIA GeForce RTX 2060 Super GPU, and an Intel Core i7-10700K CPU @ 3.80GHz.
Our evaluation framework is published in a public code repository [^4] https://github.com/Kominaru/SUB-COFI for the sake of result reproducibility. An efficient implementation of EAUC is also made available to the scientific community for use in future research in dyadic regression as a metric to evaluate model bias and fairness.
VI Results and Discussion
The results Table III includes the performance metrics (RMSE, MAE, EAUC) obtained for each model and dataset. We report results for the state-of-the-art models and the two naive Random and Dyad Average baselines.
TABLE III: Global error RMSE and MAE and eccentricity bias EAUC per model and dataset, and “fairness difficulty” $D_{KS_{\mathcal{D}}}$ per dataset.
Movielens 100K $D_{KS_{\mathcal{D}}}$: 0.415 Movielens 1M $D_{KS_{\mathcal{D}}}$: 0.412
RMSE MAE EAUC RMSE MAE EAUC
Random 1.690 ± 0.009 1.381 ± 0.011 0.416 ± 0.015 1.707 ± 0.002 1.395 ± 0.003 0.422 ± 0.011
Dyad Average 0.978 ± 0.005 0.791 ± 0.005 0.401 ± 0.003 0.958 ± 0.002 0.772 ± 0.001 0.423 ± 0.006
[…]
Glocal-K 0.882 ± 0.003 0.689 ± 0.004 0.343 ± 0.016 0.826 ± 0.003 0.643 ± 0.002 0.347 ± 0.022
GC-MC 0.891 ± 0.006 0.700 ± 0.005 0.371 ± 0.018 0.846 ± 0.005 0.665 ± 0.004 0.371 ± 0.018
MF 0.907 ± 0.008 0.711 ± 0.007 0.365 ± 0.022 0.840 ± 0.003 0.657 ± 0.003 0.366 ± 0.010
Movielens 10M $D_{KS_{\mathcal{D}}}$: 0.437 Netflix Prize $D_{KS_{\mathcal{D}}}$: 0.429
RMSE MAE EAUC RMSE MAE EAUC
Random 1.841 ± 0.001 1.503 ± 0.000 0.409 ± 0.010 1.696 ± 0.000 1.386 ± 0.000 0.432 ± 0.002
[…]
Bayesian SVD++ 0.759 ± 0.001 0.580 ± 0.001 0.328 ± 0.011
Glocal-K
[…]
Douban Monti $D_{KS_{\mathcal{D}}}$: 0.466 Tripadvisor $D_{KS_{\mathcal{D}}}$: 0.539
RMSE MAE EAUC RMSE MAE EAUC
Random 1.632 ± 0.011 1.333 ± 0.008 0.409 ± 0.019 1.826 ± 0.005 1.500 ± 0.005 0.425 ± 0.007
Dyad Average 0.771 ± 0.000 0.619 ± 0.000 0.416 ± 0.000 0.957 ± 0.004 0.737 ± 0.003 0.467 ± 0.020
Bayesian SVD++
Glocal-K 0.723 ± 0.001 0.562 ± 0.001 0.382 ± 0.002
[…]
  • In each dataset, the first group of results are naive baselines, and the second group are state-of-the-art models. Results represent performance over five runs. Other than Matrix Factorization (MF), model results include only the datasets present in their respective publications. For each dataset and metric, bold and underlined values represent the best and second best-performing model, respectively.
The results reveal that, when comparing performances between different datasets, a lower RMSE does not correlate with a lower bias (and therefore a lower EAUC). For instance, models trained on Movielens 1M showcase lower RMSE than those trained on Movielens 100K, yet exhibit equally high EAUC values and therefore similarly severe eccentricity bias. Because the optimization criterion is similar for all the models evaluated (minimizing mean errors), if we consider a single dataset we observe models with a slightly better RMSE exhibit a still high but slightly lower EAUC; in Section VI-A, we demonstrate that this slight difference between high EAUC values is not sufficient to consider the lower better model as sufficiently unbiased. Table III also contains the $D_{KS_{\mathcal{D}}}$ for each dataset: it can be observed that datasets with a higher $D_{KS_{\mathcal{D}}}$ have higher EAUC for the trained models, indicating that less uniform local distributions of observed values of users and items lead to learnt models with more severe eccentricity biases.
Figure 4 visually represents the relationship between the eccentricity of observed values and the prediction errors. Model errors are near-zero when the eccentricity is very small but increase almost linearly with eccentricity. This is particularly severe in datasets with very non-uniform observed value distributions, such as Douban Monti and Tripadvisor, in which a global optimizer will be rewarded for predicting always values near $DMV_{ui}$.
When eccentric cases hold significant importance, such as in individualized drug dosage prediction, avoiding large errors in eccentric cases is more critical for system usability and fairness than seeking near-zero errors in common, predictable scenarios. Evaluating such tasks should not be limited to global error metrics like RMSE or MAE, but also measuring entity-wise biases, which we achieve with the proposed EAUC, to promote model fairness.
Figure 4: Relationship between eccentricity of observed values and prediction error (lower is better) for each model and dataset (mean and standard deviation of five runs). For models other than MF, results are reported for datasets used in each model’s original publication.
VI-A Correlating lower EAUC and lighter eccentricity bias
As previously discussed, an unbiased model should prioritize sufficiently good predictions no matter the underlying eccentricity of the sample, rather than excelling only at predicting the $DMV_{ui}$ of the $(u,i)$ dyad -by definition, the most common test case- but making worse-than-random predictions for deviated examples with equal importance in real-world tasks. To demonstrate the ability of EAUC to measure the severity of a model’s eccentricity bias (a lower EAUC directly correlating with a less biased model, and vice-versa), we experimented with simple linear corrections as naive prediction de-biasing techniques.
For all datasets, the Train set is separated into the final Train set, used to train the dyadic regression model, and a small “Correction” set, used to compute the necessary de-biasing corrections based on the predictive behaviour observed on the trained model. In the Correction set, to uniformize the observed values distributions of users and items, we segregated the samples $(u,i,r_{ui})$ in 10x10 equally spaced bins depending on their user’s and item’s average observed values $\bar{r}_{u}$ and $\bar{r}_{i}$, and applied ML-RUS undersampling [27] to each bin to obtain the same number of samples per unique $r_{ui}$ value and bin: this avoids the learnt correction from succumbing from the same inductive bias as the original model. The correction function is then obtained by fitting on the undersampled Correction set a linear regression that takes as inputs the -possibly biased- value $\hat{r}_{ui}$ predicted by MF and the user’s and item’s average observed values $\bar{r}_{u}$ and $\bar{r}_{i}$, and must output $r_{ui}$, the real -unbiased- observed value of the sample, such that $\hat{r}_{ui_{\text{unbiased}}}=f_{corr}(\hat{r}_{ui_{\text{biased}}},\bar{r}_{u},\bar{r}_{i})$. In essence, this correction tries to learn the degree of eccentricity bias and towards which average observed value the predictions are skewed depending on the past observed values of the user and item, to appropriately de-bias Test set predictions by approximating them to their likely real value.
Table IV and Figure 5 show the effects of the corrections in the eccentricity bias and the different metrics when applied to an MF model and evaluated in each dataset’s Test set. We use two different approaches to rescale the corrected predictions back to the $[min(r_{ui}),max(r_{ui})]$ range: the Clipping approach clips the estimations to said range, while the Sigmoid approach applies a sigmoidal function and then rescales the output to the range, to avoid a collapse of corrected predictions toward the extreme values. Although they should not be interpreted as a foolproof de-biasing prediction, their effects on the prediction behaviour of MF are evident. While the corrections perform worse than the trained models for near-zero eccentricities, they maintain a much more restrained error as eccentricity increases. Importantly, this de-biasing correction exhibits a correspondence with a lower EAUC across all datasets, indicating that a conscious mitigation of the bias does signify a reduction in EAUC and vice-versa.
TABLE IV: Global error RMSE and MAE and eccentricity bias EAUC of MF and naive de-biasing techniques.
Movielens 100K $D_{KS_{\mathcal{D}}}$: 0.415 Movielens 1M $D_{KS_{\mathcal{D}}}$: 0.412
RMSE MAE EAUC RMSE MAE EAUC
Random 1.690 ± 0.009 1.381 ± 0.011 0.416 ± 0.015 1.707 ± 0.002 1.395 ± 0.003 0.422 ± 0.011
[…]
MF 0.907 ± 0.008 0.711 ± 0.007 0.365 ± 0.022 0.840 ± 0.003 0.657 ± 0.003 0.366 ± 0.010
[…]
Movielens 10M $D_{KS_{\mathcal{D}}}$: 0.437 Netflix Prize $D_{KS_{\mathcal{D}}}$: 0.429
RMSE MAE EAUC RMSE MAE EAUC
Random 1.841 ± 0.001 1.503 ± 0.000 0.409 ± 0.010 1.696 ± 0.000 1.386 ± 0.000 0.432 ± 0.002
Dyad Average 0.913 ± 0.001 0.719 ± 0.001 0.430 ± 0.007 0.964 ± 0.000 0.781 ± 0.000 0.467 ± 0.004
MF 0.782 ± 0.001 0.597 ± 0.001 0.341 ± 0.013 0.826 ± 0.001 0.635 ± 0.001 0.379 ± 0.003
[…]
MF + Linear Corr. (Sigmoid) 0,935 ± 0,008 0,741 ± 0,007 0,271 ± 0,009 0,978 ± 0,001 0,784 ± 0,001 0,289 ± 0,002
Douban Monti $D_{KS_{\mathcal{D}}}$: 0.466 Tripadvisor $D_{KS_{\mathcal{D}}}$: 0.539
RMSE MAE EAUC RMSE MAE EAUC
Random 1.632 ± 0.011 1.333 ± 0.008 0.409 ± 0.019 1.826 ± 0.005 1.500 ± 0.005 0.425 ± 0.007
[…]
MF + Linear Corr. (Sigmoid) 0,833 ± 0,014 0,660 ± 0,015 0,362 ± 0,008 1,308 ± 0,017 1,097 ± 0,015 0,372 ± 0,012
  • The two naive de-biasing techniques employ linear corrections (estimation of the real value from the biased prediction) with Clipping or Sigmoid rescaling. For each dataset and metric, bold and underlined values represent the best and second best-performing model, respectively.
Figure 5: Relationship between the eccentricity of observed values and the prediction error (lower is better) for MF and the two linear de-biasing techniques with Clipping or Sigmoid rescaling. Mean and standard deviation of five executions are reported.
VII Conclusions and Future Directions
This study provides a comprehensive examination of the impact of non-uniform local user and item observed value distributions on regression models over dyadic data. We uncover a consistent bias across state-of-the-art models where predictions tend to severely gravitate towards user and item average observed values, leading to worse-than-random errors in more eccentric samples. While RMSE and MAE, used ubiquitously as the sole evaluation metrics in the task, only measure global prediction errors, they fall short in capturing biases such as the aforementioned, posing risks of unfairness and potentially hazardous implications depending on the system’s final application.
To address this limitation, we introduce Eccentricity Area-Under the Curve (EAUC) as a complementary metric capable of evaluating model performance across all levels of eccentricity in a dataset or particular dyadic regression task. We demonstrate, in a comprehensive analysis with real-world datasets and state-of-the-art models, that a lower EAUC signifies a model that performs acceptably well for all eccentricities rather than rewarding an overemphasis on predictions near the average observed values for users and/or items. Additionally, we show that an entity-wise adaptation of the Kolmogorov-Smirnov statistic can measure a dataset’s susceptibility to induce eccentricity bias in models optimizing for global error minimization.
As we look to the future, potential avenues for further exploration involve testing these analyses on dyadic regression datasets that diverge from being surrogate tasks for learning-to-rank, especially those with greater disparities in entity-wise average observed values. While the challenges of obtaining such data are significant, it is essential for understanding the broader implications of this eccentricity-related predictive bias. Additionally, we believe future work in the field of regression over dyadic data should prioritize the development of dyadic regression models or learning optimizers less susceptible to eccentricity bias, even if their impact on global error metrics is not substantial. Such advancements would represent a crucial step toward ensuring fairness in dyadic regression models across critical applications -like biomedical or financial ones- that could influence the actions of people, companies or institutions.
Appendix A Model Hyperparameters
Table V below summarizes the hyperparameter configurations employed for various models across different datasets in the experimental setup. These configurations were either adopted from the original publications or extracted from the official repositories and source codes.
TABLE V: Hyperparameter configurations per model and dataset.
Movielens 100K Movielens 1M Movielens 10M Netflix Prize Douban Monti Tripadvisor
MF Embedding Dimension 128 512 512 512 8 128
Learning Rate 0.0005 0.001 0.001 0.001 0.001 0.0005
$L_{2}$ Weight Regularization 0.0001 0.00001 0.00001 0.00001 0.0001 0.0001
Early Stop Patience 10 10 10 10 10 10
Early Stop Delta 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
Bayesian SVD++ Embedding Dimension 10 32 128
MCMC Iterations 512 512 512
GC-MC Dropout 0.7 0.7 0.3 0.7
Hidden Layers 500, 75 500, 75 500, 75 500, 75
Learning Rate 0.01 0.01 0.01 0.01
Training Epochs 1000 3500 15000 300
Mixture Model GCN Functions 2 2 4 2
GLocal-K Auto-Encoder Hiden Layers 500, 500 500, 500 500, 500
RBF Kernel Dimensions 5, 5 5, 5 5, 5
Global Kernel Dimension 3x3 3x3 3x3
$\lambda_{2}$ $L_{2}$ Weight Regularization 20 70 10
$\lambda_{s}$ $L_{2}$ Kernel Regularization 0.006 0.018 0.022
L-BFGS-B Pre-training Iterations 5 50 5
L-BFGS-B Fine-tuning Iterations 5 10 5